# Toward Better Temporal Structures for Geopolitical Events Forecasting

### v1

@@ Abstract
Forecasting on geopolitical temporal knowledge graphs (TKGs) through the lens of large language models (LLMs) has recently gained traction. While TKGs and their generalization, hyper-relational temporal knowledge graphs (HTKGs), offer a straightforward structure to represent simple temporal relationships, they lack the expressive power to convey complex facts efficiently. One of the critical limitations of HTKGs is a lack of support for more than two primary entities in temporal facts, which commonly occur in real-world events. To address this limitation, in this work, we study a generalization of HTKGs, Hyper-Relational Temporal Knowledge Generalized Hypergraphs (HTKGHs). We first derive a formalization for HTKGHs, demonstrating their backward compatibility while supporting two complex types of facts commonly found in geopolitical incidents. Then, utilizing this formalization, we introduce the htkgh-polecat dataset, built upon the global event database POLECAT. Finally, we benchmark and analyze popular LLMs on the relation prediction task, providing insights into their adaptability and capabilities in complex forecasting scenarios111 https://github.com/usc-isi-i2/htkgh-polecat.
@@ 1 Introduction
Temporal knowledge graphs (TKGs) are one of the most common structures used to express temporal relational facts. In its simplest form, a TKG stores information as quadruples, each consisting of two entities, a relation, and a timestamp, which express a typed, directed temporal edge. Recent extensions of TKGs, such as hyper-relational temporal knowledge graphs (HTKGs), allow additional information to be added to the edge in the form of qualifiers, pairs of relations and entities (ding-etal-2024-temporal). Apart from storing information, these structures are commonly used to run prediction queries, finding the most likely element(s) of a given partial fact. One of the more challenging types of such queries is forecasting, which requires the model to make predictions in a time range it has not seen (see Section 3). Making predictions in the geopolitical domain is one of the recurrent use cases of such forecasting queries (garcia-duran-etal-2018-learning; jin-etal-2020-recurrent; 10.5555/3737916.3742366). However, a common occurrence in geopolitical events is the involvement of numerous primary entities in an event, which creates a complex fact that is outside the norms of the two-primary-entity system that HTKG relies on. In this work, we aim to address this issue through a generalization of the HTKG structure (see Figure 1 for an overview).
@@
Figure 1: An overview of how facts are expressed in TKGs, HTKGs, and HTKGHs. HTKGHs adds efficient support for complex facts involving many (i.e., more than two) primary entities, expanding the possible expression scenarios such as multi-national treaties.
@@
Examining the stored events in databases such as POLECAT (halterman2023plover), we observe that many events involve group(s) of entities interacting with each other. Specifically, two types of such events are more common: 1) Group-Type: a group of entities (>2) having the same mutual relationship, and 2) Set2Set-Type: a group of entities taking an action against another group of entities (>2 total entities). However, trying to express such events within existing structures is difficult, as the simple structure of directed edges between two entities fails to properly represent them without imposing further issues, such as sparsity (see Section 3.4). To mitigate this problem, this work derives a formalization for hyper-relational temporal knowledge generalized hypergraphs (HTKGHs), which support an arbitrary number of primary entities and second-order edges, enabling efficient expression of more complex geopolitical events (see Section 4). Moreover, we introduce a new dataset, htkgh-polecat, which utilizes the HTKGH formalization, based on the POLECAT database. Our statistics show that roughly one in four facts in htkgh-polecat follows one of the newly supported complex formats (see Section 5).
@@
In recent years, LLMs have been extensively studied as a new approach to solving TKG queries, as they offer an easy-to-use approach that works out of the box with many structures via small amounts of prompting (lee-etal-2023-temporal; xia-etal-2024-chain). Such capabilities mainly stem from the extensive pre-training stage they undergo, where they are exposed to copious amounts of information, learning complex syntactical and semantic relationships between concepts and entities. Notably, these capabilities extend beyond simple memorization, as they have demonstrated some level of general pattern recognition even when both semantic and syntactic information are hidden (lee-etal-2023-temporal). One of the common tasks on TKGs is relation prediction, which in many ways mirrors common logical reasoning tasks for LLMs. In this work, we evaluate the reasoning capabilities of LLMs over complex facts using the newly introduced dataset. Our experiments highlight the impressive adaptiveness of LLMs when reasoning over information that is hidden or tampered with in a way that goes against the models’ prior beliefs (see Section 6).
@@
  • We analyze popular LLMs on the relation prediction task, showcasing their high performance and resilience, while highlighting the effects of various related confounding factors.
@@ 2 Related Work
TKGs are known to struggle with higher-order relational details (lu2025two). Within the TKG structure, reification (noy2006defining) represents n-ary facts with an intermediate node connected to several binary predicates. chebba2018attributed proposes attributed relations and modeling n-ary relations with blank nodes. However, these approaches unnecessarily increase the graph size, leading to extreme inefficiencies. Beyond the TKG structure, HTKGs (ding-etal-2024-temporal) allow contextual attributes by adding qualifiers, disambiguating between otherwise identical TKG quadruples. N-TKGs (hou-etal-2023-temporal) extend TKGs’ quadruples into n-tuples, where relations connect $n$ entities with assigned roles, allowing for n-ary facts. UniHR (liu2024unihr) introduces a hierarchical data representation module (HiDR) to jointly model heterogeneous information. However, these extensions fail to move beyond the two-primary-entity system.
@@
Traditional methods for TKG forecasting include time-aware embedding models (goel2020diachronic; han2020xerte), temporal random walk-based approaches (sun2021timetraveler; liu2022tlogic), and graph-based approaches (zhu2021learning; li2021temporal; xu2023temporal). More recently, a wave of LLM-based approaches has emerged, leveraging the pattern recognition capabilities of LLMs. lee-etal-2023-temporal studies LLMs’ ability to perform TKG forecasting based on in-context learning. GenTKG (liao2023gentkg) incorporates retrieval-augmented generation (RAG) into their method. zrLLM (ding2023zrllm) relies on LLMs’ semantic reasoning abilities to model unseen relations. Nonetheless, LLMs’ reasoning capabilities on complex facts remain understudied.
@@
Geopolitical event databases are rich sources for building TKGs, offering chronologically ordered facts involving several actors in one or more actions (e.g., wars, treaties, etc.). Previously, snapshots of GDELT (leetaru2013gdelt) and ICEWS (DVN/28075_2015), both part of broad initiatives to catalog global human behavior and predict conflicts, were used as TKG reasoning benchmarks. Other examples are ACLED (raleigh2010introducing; semnani-etal-2025-lemonade), containing location and battle event information related to eight conflict countries in West and Central Africa from 1960 to the present, and UCDP (sundberg2012introducing; sundberg2013introducing), containing global data on non-state conflicts from 1989 to 2024. In this work, we utilize POLECAT (halterman2023plover; halterman2023creating), which stores cooperative and hostile geopolitical events from 2018 to 2024.
@@
Figure 2: (a) The original temporal fact representing a joint trade deal between $\{\text{China, Japan, and South Korea}\}$ on $\{\text{Cars, Chips, and Oil}\}$. (b) The decomposed version of the original fact designed for HTKGs. Notably, the decomposed version requires a lot of redundancy to represent, and without an additional entity representing the group of countries, it is indistinguishable from three separate trade deals between these countries.
@@ 3.1 Definitions
Following the definition by ding-etal-2024-temporal, we define an HTKG $\mathcal{H}$ as an attributed TKG where each quadruple is accompanied by a set of tuples containing additional information about the temporal fact. Formally, let $\mathcal{G}$ be a TKG containing the primary quadruples, we define $\mathcal{H}$ as
@@ 4 Hyper-Relational Temporal Knowledge Generalized Hypergraph (HTKGH)
Unidirectional vs. Bidirectional Edges
$(\{v\},r_{u},\{u\},t),Q)\;\;\text{vs.}\;\;(\{v,u\},r_{b},\{\},t),Q)$
@@ 5 Dataset
POLitical Event Classification, Attributes, and Types (POLECAT) (halterman2023plover) is a new global events database built to succeed the Integrated Conflict Early Warning System (ICEWS) (DVN/28075_2015). POLECAT utilizes the Political Language Ontology for Verifiable Event Records (PLOVER) codebook for coding event data, a new event-mode-context ontology designed to be more general, easy to implement, and extendable, replacing the Conflict and Mediation Event Observations (CAMEO) framework (gerner2002conflict) utilized in ICEWS. Compared to CAMEO, PLOVER defines 18 event types, aggregating over many of the more than 250 CAMEO codes. To create htkgh-polecat, we utilize the POLECAT data from Jan 2018 to Jul 2024. Appendix C provides more details on dataset construction and statistics, such as entity and relation composition.
@@
Regular True True True 49.6 34.5 48.7 42.3 58.1 55.4 49.0 60.6 25.5 64.7 63.1 63.9
True True False 43.3 28.9 23.8 35.4 44.4 40.4 39.9 45.9 15.4 50.0 48.3 47.2
True False True 45.8 28.5 32.9 40.1 50.5 47.9 43.4 52.0 18.8 56.1 54.8 54.1
True False False 41.2 25.4 6.6 28.1 36.9 32.3 34.5 37.8 11.5 40.7 39.8 36.4
False True True 38.5 21.4 22.8 34.8 40.8 36.7 35.6 44.3 12.4 48.4 44.8 43.0
False True False 30.5 11.9 4.0 26.4 30.1 23.7 26.5 32.2 8.3 35.5 31.4 29.2
False False True 35.6 14.8 4.7 25.8 30.2 27.0 25.5 32.8 7.9 37.7 34.3 30.3
False False False 27.3 5.0 0.0 25.5 28.4 21.3 22.0 29.4 6.6 34.2 28.8 27.1
@@ 6 Experiments
In these experiments, we focus on the relation prediction task, which is similar to the logical reasoning tasks, a desirable ability for LLMs.
(a) Gemma-3
(b) Qwen3
Figure 5: Relation prediction accuracy (%) on htkgh-polecat for different model sizes in the (a) Gemma-3 and (b) Qwen3 families. Legend: N $\rightarrow$ non-thinking, T $\rightarrow$ thinking.
@@ 6.1 Evaluation Setup
We select 13 models from common publishers such as Google and Qwen, based on size and type, enabling further analysis on these critical factors (see Appendix H for more details). For size, our models range from 270M to 20B. For type, we focus on two common types: non-thinking (i.e., instruction-tuned) and thinking. Finally, to better analyze their performance, we allow thinking models to generate up to 16384 tokens while non-thinking models are limited to 14 tokens (i.e., max length of the relations, nine, plus five), all using their preferred decoding strategy and parameters.
@@ 6.2 Experimental Results
To better understand LLMs’ performance, we compare them against three baselines that use simple heuristics that have been shown to either impose strong bias in LLMs lee-etal-2023-temporal or have strong performance in TKG forecasting benchmarks 10.24963/ijcai.2024/444. More specifically, given the historical context $\mathcal{H}$, we calculate 1) Frequency as the relation with the most occurrences, 2) Recency as the relation that comes last, and 3) Copy as the last relation that appears with the same elements as the query. Table 1 (Regular variation) presents our experimental results on nine different state-of-the-art non-thinking and thinking LLMs. Looking at the non-thinking category, only gemma-3-12b-it and Qwen3-4B-Instruct-2507 relatively consistently beat the heuristics, showcasing critical improvements with size increases (gemma-3-4b-it vs. gemma-3-12b-it) and alignment improvements [^4] Based on the release card, it seems that Qwen3-*-2507 models have gone through an improved post-training phase (vs. Qwen3-* models) rather than a full pre-training phase. (Qwen3-8B (non-thinking) vs. Qwen3-4B-Instruct-2507). Moreover, we see that with more rigorous filtering, which leads to more relevant facts being present in the history, LLMs tend to improve faster than the heuristics (i.e., increasing their lead margin or closing the performance gap), showcasing the potential gains of improving history retrievers, similar to the findings of xia-etal-2024-chain. In the thinking category, we observe more cases where the models beat the heuristics, along with similar observations on the filtering; however, we notice two surprising results: 1) DeepSeek-R1-Distill-Qwen-7B does not perform well at all, which could partially be explained by misformatting rates (see Appendix G), and 2) Qwen-3 models beat the significantly larger gpt-oss model by substantial margins in many cases, showcasing the power of smaller well-trained models. Moreover, we note a massive jump of +6.0-7.6% from non-thinking to thinking variation of the Qwen3 models, which highlights the effectiveness of test-time scaling for the relation prediction task on TKGs.
Variation Filters Heuristics GNN LLM
[…]
Regular True True True 51.7 39.1 56.0 48.2 49.1 48.8 49.1 48.9 63.3 59.4 52.0 65.5 20.6 70.2 65.8 69.3
True True False 47.0 33.9 33.3 48.7 48.9 49.4 49.0 41.9 50.5 46.2 44.2 51.5 13.8 57.1 53.8 55.5
True False True 49.1 31.2 41.2 48.5 48.6 48.2 48.7 46.8 56.8 54.2 47.7 58.4 18.1 62.3 60.0 61.2
True False False 44.9 29.0 12.8 48.2 49.4 48.2 48.2 33.6 43.5 39.7 39.3 43.7 11.8 47.7 45.3 43.7
False True True 43.0 25.5 28.5 48.2 49.1 49.3 48.7 39.6 46.4 41.2 40.5 49.0 12.0 53.0 48.4 49.5
False True False 33.7 14.7 8.6 49.1 48.4 48.9 48.8 30.9 35.8 27.6 30.5 37.0 7.3 40.4 36.0 35.1
False False True 39.1 17.3 7.3 49.1 49.2 49.5 49.1 29.4 34.5 30.7 30.1 37.6 7.3 41.2 37.7 33.9
False False False 26.4 4.3 4.8 49.8 49.4 49.2 49.7 28.2 32.3 23.8 23.8 31.9 4.9 37.9 30.5 30.0
@@
While the observed improvements using test-time scaling are a sign of LLMs’ going beyond memorization, we further experiment with two shuffled variations of our dataset, where 1) only countries are shuffled, and 2) countries and relations are shuffled. Table 1 (Shuffled variations) presents our experimental results for this experiment. In the entity-only variation, we see that generally non-thinking models’ accuracy slightly increases (up to +5.3%) while thinking models’ accuracy slightly decreases (up to -1.9%), which showcases their resilience and context-dependence to a great degree, similar to the findings of lee-etal-2023-temporal. However, when we also shuffle the relations, the behavior of many of the models becomes erratic, with some gaining up to +9.6% accuracy and some losing up to -11.3% accuracy. This phenomenon highlights the importance of testing LLMs under extreme settings where the provided information goes against the model’s beliefs to different degrees.
Q3. What is the effect of the number of contextual samples?
One of the critical considerations of using LLMs for TKG-related tasks is the number of facts that could be included in the context. Figure 4 illustrates our experimental results with respect to the number of contextual samples. As is evident, most models benefit from the inclusion of more context, DeepSeek-R1-Distill-Qwen-7B being the only model that degrades with more samples. Moreover, we observe that scenarios with stricter filtering benefit more from an increase in samples. These results illustrate a promising trend in context scalings for more complex TKG-related predictions. Note that some models allow us to scale the samples further, but due to resource constraints, we leave these experiments to future work.
Q4. What is the effect of the model size?
While many training factors are often undisclosed, model size has always been one of the consistent predictors of performance. Figure 5 showcases a comparison with regard to model size within the Gemma-3 and Qwen3 families (see Appendix J for more filtering variations). In both families, the performance generally increases as model sizes increase, especially past the 1B size threshold. The most significant exceptions are the Qwen3-4B-*-2507 models, which even beat the much larger Qwen3-14B model, underscoring the critical nature of other factors that must be studied for model selection.
Q5. How good are LLMs compared to supervised graph-based models?
We experiment with two graph neural network (GNN) models, both based on first encoding fact-level information, then aggregating facts within a time window with either a permutation invariant bagging or a more expressive hypergraph approach, and finally using a transformer-based encoder to capture temporal trends (see Appendix K for more granular details on the GNN-based models). Since these supervised models need a training phase, we split the dataset into a train (i.e., 2018-2022) and a test (i.e., 2023-2024) set. Then, we randomly sample and hold out 10% of the training data as a validation set. Finally, we calculate the performance of the model on the validation set after each epoch and select the version with the best validation result for testing.
@@
Table 2 presents our experimental results comparing GNN-based models with our baselines and selected LLMs. As is evident, in scenarios where we have weak filtering over contextual information, GNNs consistently beat LLMs, which we can potentially attribute to the information stored in embeddings and learned weights. However, as we tighten our filtering, LLMs start to gain momentum, beating GNN-based models, whose performance stays flat, by as much as 21%. These results emphasize the importance of retrieving high-quality context for predictions again. Regarding our GNN-based models, we hypothesize that the current window encoding formulation introduces an information bottleneck by collapsing all fact representations, which leads to their flat performance across different filtered settings. Given the numerous different modeling choices and designs for GNNs, we leave further investigations to future work.

### v2

@@ Abstract
Forecasting on geopolitical temporal knowledge graphs (TKGs) through the lens of large language models (LLMs) has recently gained traction. While TKGs and their generalization, hyper-relational temporal knowledge graphs (HTKGs), offer a straightforward structure to represent simple temporal relationships, they lack the expressive power to convey complex facts efficiently. One of the critical limitations of HTKGs is a lack of support for more than two primary entities in temporal facts, which commonly occur in real-world events. To address this limitation, in this work, we study a generalization of HTKGs, Hyper-Relational Temporal Knowledge Generalized Hypergraphs (HTKGHs). We first derive a formalization for HTKGHs, demonstrating their backward compatibility while supporting two complex types of facts commonly found in geopolitical incidents. Then, utilizing this formalization, we introduce the htkgh-polecat dataset, built upon the global event database POLECAT. Finally, we benchmark and analyze popular LLMs on our dataset, providing insights into 1) the positive impact of utilizing the HTKGH formalization compared to existing ones and 2) LLMs’ adaptability and capabilities in complex forecasting tasks111 https://github.com/usc-isi-i2/htkgh-polecat.
@@ 1 Introduction
Temporal knowledge graphs (TKGs) are one of the most common structures used to express temporal relational facts. In its simplest form, a TKG stores information as quadruples, each consisting of two entities, a relation, and a timestamp, which express a typed, directed temporal edge. Recent extensions of TKGs, such as hyper-relational temporal knowledge graphs (HTKGs), allow additional information to be added to the edge in the form of qualifiers, pairs of relations and entities (Ding et al., 2024). Apart from storing information, these structures are commonly used to run prediction queries, finding the most likely element(s) of a given partial fact. One of the more challenging types of such queries is forecasting, which requires the model to make predictions in a time range it has not seen (see Section 3). Making predictions in the geopolitical domain is one of the recurrent use cases of such forecasting queries (García-Durán et al., 2018; Jin et al., 2020; Gastinger et al., 2025). However, a common occurrence in geopolitical events is the involvement of numerous primary entities in an event, which creates a complex fact that is outside the norms of the two-primary-entity system that HTKG relies on. In this work, we aim to address this issue through a generalization of the HTKG structure (see Figure 1 for an overview).
@@
Figure 1: An overview of how facts are expressed in TKGs, HTKGs, and HTKGHs. HTKGHs add efficient support for complex facts involving many (i.e., more than two) primary entities, expanding the possible expressions to scenarios such as multi-national treaties.
@@
Examining the stored events in databases such as POLECAT (Halterman et al., 2023a), we observe that many events involve group(s) of entities interacting with each other. Specifically, two types of such events are more common: 1) Group-Type: a group of entities (>2) having the same mutual relationship, and 2) Set2Set-Type: a group of entities taking an action against another group of entities (>2 total entities). However, trying to express such events within existing structures is difficult, as the simple structure of directed edges between two entities fails to properly represent them without imposing further issues, such as sparsity (see Section 3.4). To mitigate this problem, this work derives a formalization for hyper-relational temporal knowledge generalized hypergraphs (HTKGHs), which support an arbitrary number of primary entities and second-order edges, enabling efficient expression of more complex geopolitical events (see Section 4). Moreover, we introduce a new dataset, htkgh-polecat, which utilizes the HTKGH formalization, based on the POLECAT database. Our statistics show that roughly one in four facts in htkgh-polecat follows one of the newly supported complex formats (see Section 5).
@@
In recent years, LLMs have been extensively studied as a new approach to solving TKG queries, as they offer an easy-to-use approach that works out of the box with many structures via small amounts of prompting (Lee et al., 2023; Xia et al., 2024). Such capabilities mainly stem from the extensive pre-training stage they undergo, where they are exposed to copious amounts of information, learning complex syntactical and semantic relationships between concepts and entities. Notably, these capabilities extend beyond simple memorization, as they have demonstrated some level of general pattern recognition even when both semantic and syntactic information are hidden (Lee et al., 2023). In this work, we evaluate the reasoning capabilities of LLMs over complex facts using the newly introduced dataset. Our experiments highlight the impressive adaptiveness of LLMs when reasoning over information that is hidden or tampered with in a way that goes against the models’ prior beliefs (see Section 6). Moreover, we demonstrate the empirical performance and inference time improvements that can be achieved using HTKGH over the existing structures.
@@
  • We adapt and analyze popular LLMs and graph-based models on our dataset, showcasing critical gains that can be achieved using the HTKGH formalization on performance and inference time, while also highlighting the impact of various confounding factors.
@@ 2 Related Work
TKGs are known to struggle with higher-order relational details (Lu et al., 2025). Within the TKG structure, reification (Noy et al., 2006) represents n-ary facts with an intermediate node connected to several binary predicates. Chebba et al. (2018) proposes attributed relations and modeling n-ary relations with blank nodes. However, these approaches unnecessarily increase the graph size, leading to extreme inefficiencies. Beyond the TKG structure, HTKGs (Ding et al., 2024) allow contextual attributes by adding qualifiers, disambiguating between otherwise identical TKG quadruples. N-TKGs (Hou et al., 2023) extend TKGs’ quadruples into n-tuples, where relations connect $n$ entities with assigned roles, allowing for n-ary facts. UniHR (Liu et al., 2024) introduces a hierarchical data representation module (HiDR) to model heterogeneous information jointly. However, these extensions fail to move beyond the two-primary-entity system.
@@
Traditional methods for TKG forecasting include time-aware embedding models (Goel et al., 2020; Han et al., 2020), temporal random walk-based approaches (Sun et al., 2021; Liu et al., 2022), and graph-based approaches (Zhu et al., 2021; Li et al., 2021; Xu et al., 2023). More recently, a wave of LLM-based approaches has emerged, leveraging the pattern recognition capabilities of LLMs. Lee et al. (2023) studies LLMs’ ability to perform TKG forecasting based on in-context learning. GenTKG (Liao et al., 2023) incorporates retrieval-augmented generation (RAG) into their method. zrLLM (Ding et al., 2023) relies on LLMs’ semantic reasoning abilities to model unseen relations. Nonetheless, LLMs’ reasoning capabilities on complex facts remain understudied.
@@
Geopolitical event databases are rich sources for building TKGs, offering chronologically ordered facts involving several actors in one or more actions (e.g., wars, treaties, etc.). Previously, snapshots of GDELT (Leetaru and Schrodt, 2013) and ICEWS (Boschee et al., 2015), both part of broad initiatives to catalog global human behavior and predict conflicts, were used as TKG reasoning benchmarks. Other examples are ACLED (Raleigh et al., 2010; Semnani et al., 2025), containing location and battle event information related to eight conflict countries in West and Central Africa from 1960 to the present, and UCDP (Sundberg et al., 2012; Sundberg and Melander, 2013), containing global data on non-state conflicts from 1989 to 2024. In this work, we utilize POLECAT (Halterman et al., 2023a, b), which stores cooperative and hostile geopolitical events from 2018 to 2024.
@@
Figure 2: (a) The original temporal fact representing a joint trade deal between $\{\text{China, Japan, and South Korea}\}$ on $\{\text{Cars, Chips, and Oil}\}$. (b) The decomposed version of the original fact designed for HTKGs. Notably, the decomposed version requires a lot of redundancy to represent. Moreover, without an additional entity representing the group of countries, it is indistinguishable from three separate trade deals between these countries.
@@ 3.1 Definitions
Following the definition by Ding et al. (2024), we define an HTKG $\mathcal{H}$ as an attributed TKG where each quadruple is accompanied by a set of tuples containing additional information about the temporal fact. Formally, let $\mathcal{G}$ be a TKG containing the primary quadruples, we define $\mathcal{H}$ as
@@ 4 Hyper-Relational Temporal Knowledge Generalized Hypergraph (HTKGH)
The derived HTKGH formalization can be used by both graph neural networks (GNNs) and LLMs approaches. While our focus is mostly on studying LLMs, specifically how they perform on these highly complex structures, Section 6 and Appendix M provide a series of experiments and mechanisms to adapt GNNs.
@@ 5 Dataset
POLitical Event Classification, Attributes, and Types (POLECAT) (Halterman et al., 2023a) is a new global events database built to succeed the Integrated Conflict Early Warning System (ICEWS) (Boschee et al., 2015). POLECAT utilizes the Political Language Ontology for Verifiable Event Records (PLOVER) codebook for coding event data, a new event-mode-context ontology designed to be more general, easy to implement, and extendable, replacing the Conflict and Mediation Event Observations (CAMEO) framework (Gerner et al., 2002) utilized in ICEWS. Compared to CAMEO, PLOVER defines 18 event types, aggregating over many of the more than 250 CAMEO codes. To create htkgh-polecat, we utilize the POLECAT data from Jan 2018 to Jul 2024. Appendix C provides more details on dataset construction and statistics, such as entity and relation composition.
@@
Regular True True True 40.2 29.4 19.4 19.9 33.8 28.8 28.0 36.8 13.7 36.7 35.9 34.8
True True False 42.5 28.9 15.4 24.5 35.9 29.9 31.5 39.1 12.7 39.1 38.9 34.8
True False True 42.9 30.3 16.6 23.2 36.3 30.8 30.8 39.6 13.1 39.0 38.3 36.1
True False False 43.3 27.3 12.6 24.3 36.3 30.9 32.7 40.0 12.2 40.6 38.9 34.7
False True True 36.4 22.8 10.6 19.7 29.5 26.4 27.5 34.2 9.7 34.7 34.8 30.3
False True False 31.5 19.8 6.7 20.2 27.3 24.1 28.0 32.6 7.5 31.9 31.3 28.2
False False True 34.8 19.6 6.4 19.1 27.8 25.9 28.0 33.3 8.3 34.3 33.2 28.6
False False False 26.5 15.0 2.5 16.0 23.4 20.9 26.7 29.3 6.8 29.9 29.3 24.6
@@ 6 Experiments
In these experiments, we mostly analyze the relation prediction task, which is similar to the common logical reasoning tasks for LLMs.
@@ 6.1 Evaluation Setup
Variation Filters Heuristics GNN LLM
[…]
Regular True True True 46.7 34.7 21.7 48.2 49.1 48.8 49.1 23.0 38.2 32.4 31.4 40.6 12.2 41.0 39.7 39.3
True True False 46.4 31.2 17.4 48.7 48.9 49.4 49.0 26.8 38.9 32.4 33.8 42.2 11.7 41.7 42.0 38.0
True False True 46.9 33.1 18.6 48.5 48.6 48.2 48.7 25.4 39.5 34.3 33.6 42.8 12.2 42.8 41.5 39.9
True False False 46.2 28.8 14.2 48.2 49.4 48.2 48.2 26.7 38.7 33.9 34.8 43.0 12.2 43.4 41.4 38.6
False True True 40.8 26.0 12.4 48.2 49.1 49.3 48.7 22.0 33.6 29.6 30.4 37.1 9.3 38.2 37.4 33.2
False True False 33.3 21.0 8.0 49.1 48.4 48.9 48.8 23.1 30.2 26.9 31.0 35.3 8.0 34.9 35.0 31.2
False False True 37.3 22.0 7.6 49.1 49.2 49.5 49.1 21.8 31.2 28.6 29.6 36.6 8.1 36.6 35.5 30.6
False False False 25.0 15.0 2.4 49.8 49.4 49.2 49.7 18.8 26.1 23.8 28.5 32.5 6.4 32.8 31.1 26.9
@@
We select 9 models from common publishers such as Google and Qwen, based on size and type, enabling further analysis on these critical factors (see Appendix G for more details). For size, our models range from 4B to 20B. For type, we focus on two common types: non-thinking (i.e., instruction-tuned) and thinking. Finally, to better analyze their performance, we allow thinking models to generate up to 16384 tokens while non-thinking models are limited to 14 tokens (i.e., max length of the relations, nine, plus five), all using their preferred decoding strategy and parameters.
@@ 6.2 Experimental Results
To better understand LLMs’ performance, we compare them against three baselines that use simple heuristics that have been shown to either impose strong bias in LLMs (Lee et al., 2023) or have strong performance in TKG forecasting benchmarks Gastinger et al. (2024). More specifically, given the historical context $\mathcal{H}$, we calculate 1) Frequency as the relation with the most occurrences, 2) Recency as the relation that comes last, and 3) Copy as the existence of a contextual sample with the same non-temporal elements as the test sample. Table 1 (Regular variation) presents our experimental results on nine different state-of-the-art non-thinking and thinking LLMs. As is evident, LLMs are only able to beat the heuristics in very few scenarios, which showcases the difficulty of the task and potential for future improvements. Looking at the non-thinking category, gemma-3-12b-it and Qwen3-4B-Instruct-2507 consistently beat their peers in their family, showcasing critical improvements with size increases (gemma-3-4b-it vs. gemma-3-12b-it) and alignment improvements [^4] Based on the release card, it seems that Qwen3-*-2507 models have gone through an improved post-training phase (vs. Qwen3-* models) rather than a full pre-training phase. (Qwen3-8B (non-thinking) vs. Qwen3-4B-Instruct-2507). Moreover, we see that with entity filtering, which leads to more relevant facts being present in the history, all models tend to improve, showcasing the potential gains of improving history retrievers, similar to the findings of Xia et al. (2024). In the thinking category, we notice similar observations on the filtering; however, we notice two surprising results: 1) DeepSeek-R1-Distill-Qwen-7B does not perform well at all, which could partially be explained by misformatting rates (see Appendix H), and 2) Qwen-3 models beat the significantly larger gpt-oss model by substantial margins in many cases, showcasing the power of smaller well-trained models. Moreover, we note a massive jump of +2.7-8.4% from non-thinking to thinking variation of the Qwen3 models, which highlights the effectiveness of test-time scaling for the relation prediction task on TKGs.
Figure 4: Relation prediction accuracy (%) and inference time comparison on Qwen3-4B-Thinking-2507 between HTKGH and reified HTKG.
@@
While the observed improvements using test-time scaling are a sign of LLMs’ going beyond memorization, we further experiment with two shuffled variations of our dataset, where 1) only countries are shuffled, and 2) countries and relations are shuffled. Table 1 (Shuffled variations) presents our experimental results for this experiment. In the entity-only variation, we see that generally non-thinking models’ accuracy slightly increases (up to +1.4%) while thinking models’ accuracy slightly decreases (up to -2.3%), which showcases their resilience and context-dependence to a great degree, similar to the findings of Lee et al. (2023). However, when we also shuffle the relations, the behavior of many of the models becomes erratic, with some gaining up to +9.2% accuracy and some losing up to -6.8% accuracy. This phenomenon highlights the importance of testing LLMs under extreme settings where the provided information goes against the model’s beliefs to different degrees.
(a) gemma-3-12b-it
(b) Qwen3-4B-Thinking-2507
Figure 5: Link prediction performance (i.e., Jaccard Index) and inference time on htkgh-polecat between HTKGH and reified HTKG using the best non-thinking and thinking models, respectively.
Q3. How good are LLMs compared to supervised graph-based models?
@@
We experiment with two GNN-based models, both based on first encoding fact-level information, then aggregating facts within a time window with either a permutation invariant bagging or a more expressive hypergraph approach, and finally using a transformer-based encoder to capture temporal trends (see Appendix M for more granular details on the GNN-based models). Since these supervised models need a training phase, we split the dataset into a train (i.e., 2018-2022) and a test (i.e., 2023-2024) set. Then, we randomly sample and hold out 10% of the training data as a validation set. Finally, we calculate the performance of the model on the validation set after each epoch and select the version with the best validation result for testing. Table 2 presents our experimental results comparing GNN-based models with our baselines and selected LLMs. As is evident, in all scenarios, GNNs consistently beat LLMs, which we can potentially attribute to the information stored in embeddings and learned weights. However, as we tighten our filtering, LLMs start to gain momentum, reducing the gap to supervised GNN-based models, whose performance stays flat, and getting as close as -6%. These results emphasize the importance of retrieving high-quality context for predictions again. Regarding our GNN-based models, we hypothesize that the current window encoding formulation introduces an information bottleneck by collapsing all fact representations, which leads to their flat performance across different filtered settings. Given the numerous different modeling choices and designs for GNNs, we leave further investigations to future work.
Q4. How does HTKGH compare to HTKG?
To better understand the direct impact of the new structure on the same underlying data, we compare HTKGH to reified HTKG using htkgh-polecat. To this end, we reify each contextual sample by first adding a new node and then connecting each element to that node using a role-specific relation (e.g., actor, relation, etc.). Moreover, for each query, we follow the same procedure but leave the relation edge empty so that the model can predict it. This approach allows us to provide the model with the same amount of information as HTKGH samples during the test time, making the comparison fair. Figure 4 presents a comparison between HTKGH and reified HTKG on relation prediction accuracy and inference time. As is evident, HTKGH beats reified HTKG by +4.7% on performance and reduces inference time by $\sim$56.8%. These results showcase the significant impact of utilizing HTKGH compared to more traditional common structures. Note that we were not able to run decomposed HTKG (Ding et al., 2024), as our models ran out of context length (i.e., 32768 tokens) for 100 contextual samples.
Q5. How does HTKGH impact link prediction performance?
While most of our experiments have been focused on relation prediction, to expand our insights into the impact of HTKGH, we compare it to reified HTKG on the link prediction task. Given the structural shift from two-entity structures to set-based structures, we opt to use 1) a set-based prediction method (i.e., asking the model to predict a set of entities instead of one entity) and 2) the Jaccard Index instead of traditional single-entity recall-based metrics such as hits at $k$. Figure 5 presents a comparison between HTKGH and reified HTKG on link prediction performance and inference time using gemma-3-12b-it and Qwen3-4B-Thinking-2507, the best non-thinking and thinking models, respectively. As is evident, similar to our previous findings, beats reified HTKG by as much as 21.29% on performance and reduces inference time by $\sim$57.3%. These results once more highlight the critical benefits of utilizing HTKGH compared to existing common structures.